OpenStudy (anonymous):

The ordered pairs below represent a relation between x and y (-3,-3)(-2,0)(-1,3)(0,6)(1,9)(2,12) Could this be a linear function and why?!

OpenStudy (anonymous):

you could choose two of the points and write the equation of the line joining those points. Then plug in the other points and see if they make the sentence true.

OpenStudy (anonymous):

The easy way to figure it out, is by graphing the coordinates.

OpenStudy (anonymous):

Let's choose (1,9) and (2,12) The slope if 12-9/2-1 or 3

OpenStudy (anonymous):

y-9=3(x-1) y=3x+6

OpenStudy (anonymous):

Now plug in the other points and see if they are on that line.

OpenStudy (anonymous):

They all work. so yes, those points are collinear. Because they all make the sentence y=3x+6 true

OpenStudy (anonymous):

Yes, but would it be because of the distance between consecutive x values are constant ?

OpenStudy (anonymous):

Yes. Correct!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Dont forget to medal!